OpenStudy (anonymous):

\[3\sqrt{4x-2+1}=\sqrt{36+1}\]

OpenStudy (anonymous):

Lets start by squaring both sides. This will remove the radicals. \[9(4x-2+1)=37\] And now try dividing both sides by 9: \[4x-1=37/9\]

OpenStudy (anonymous):

i wouldnt know thats how the question is

OpenStudy (anonymous):

is the answer 11/4?

OpenStudy (loser66):

23/18

OpenStudy (anonymous):

thats not one of my options

OpenStudy (anonymous):